OpenStudy (hexagon001):

what is the inverse of this taylor series about 0? [all in the context of complex analysis]

OpenStudy (hexagon001):

\[z+z^2-\frac{ z^3 }{ 2! }+\frac{ z^5 }{ 4! }-\frac{ z^7 }{ 6! }+...\]

OpenStudy (hexagon001):

im getting.. f^-1(w)=w-w^2+0.5w^3